# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ Acknowledgements
$\Phi_{b;(k_{1},\ldots,k_{r};s)}(x):=\sum_{i_{1},\ldots,i_{r}\ \mathrm{distinct}}\Big(\prod_{j=1}^{r}x_{i_{j}}^{k_{j}}\Big)\,T^{(r)}_{r,s}(x_{i_{1}},\ldots,x_{i_{r}}),$
the probability that $b$ blocks thrown uniformly on $[0,1]$ with a paintbox $x$ form a partition of type $k_{1},\ldots,k_{r};s$. Here $T^{(r)}_{r,s}$ is the inclusion–exclusion term enforcing that the $s$ singletons fall into distinct intervals from [10, p. 41, Appendix A]. Then, conditional on $x$, we have
$\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)=\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)+\Delta_{N}(x),\qquad\text{with }|\Delta_{N}(x)|\leq\frac{K_{b}}{N}.$
That $|\Delta_{N}(x)|$ is $O(\frac{1}{N})$ follows simply from the observation that the probability of the $n\ll N\sum_{i}x_{i}$ sample lineages falling into each of the intervals of $x$ are asymptotically independent, with corrections at worst of order $\frac{1}{N}$, which can be uniformly controlled by compactness of $\Delta$. Therefore,
$\frac{1}{c_{N}}\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)\right]=\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)\right]+o(1),$
since the $\frac{1}{N}$ coupling error is multiplied by $\Phi_{N}(A_{\varepsilon})/c_{N}=O_{\varepsilon}(1)$. Here $O_{\varepsilon}(1)$ simply means that, the term is $O(1)$ for any fixed $\varepsilon>0$. By Assumption (Appendix A.1), vague convergence on $\Delta\setminus\{\mathbf{0}\}$ yields
$\lim_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)\right]=2\int_{A_{\varepsilon}}\frac{\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
@@
$2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
which coincides with the transition for $\Xi\!\restriction_{\Delta\setminus\{\mathbf{0}\}}=2\Xi^{\prime}$.
For $x\in A_{\varepsilon}^{\mathrm{c}}$ we have $\langle x,x\rangle\leq\varepsilon$, hence $\max_{i}x_{i}\leq\|x\|_{2}\leq\varepsilon^{1/2}$. Then
@@
Consequently,
$\mathbb{P}(\text{non-binary}\mid x)\;\leq\;C_{b}\!\left(\sum_{i}x_{i}^{3}+\Big(\sum_{i}x_{i}^{2}\Big)^{2}\right)\;\leq\;C_{b}\big(\varepsilon^{1/2}+\varepsilon\big)\sum_{i}x_{i}^{2}\;\leq\;2C_{b}\,\varepsilon^{1/2}\sum_{i}x_{i}^{2}$
for some constant $C_{b}$ depending only on $b$. Therefore, using $\lim_{N\to\infty}\frac{1}{c_{N}}\mathbb{E}\!\big[\sum_{i}x_{i}^{2}\big]=2\Xi^{\prime}(\Delta\setminus{\{\mathbf{0}}\})$,
$\limsup_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(\text{non-binary}\mid x)\big]\;\leq\;2C_{b}\,\varepsilon^{1/2}\,\limsup_{N\to\infty}\frac{1}{c_{N}}\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]\;=\;4C_{b}\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})\,\varepsilon^{1/2}.$
@@
$\lim_{\varepsilon\downarrow 0}\;\limsup_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(\text{non-binary}\mid x)\big]=0.$
@@
and on $A_{\varepsilon}^{\mathrm{c}}$ the error satisfies
@@
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]=\frac{1}{c_{N}}\,\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]-\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}}\sum_{i}x_{i}^{2}\big].$
By the normalization and the vague convergence assumption (Appendix A.1),
@@
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]\;\to\;2\,\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})-2\,\Xi^{\prime}(A_{\varepsilon})\xrightarrow[\varepsilon\downarrow 0]{}\;0.$
By the normalization so that all binary mergers occur at rate $2c_{N}$, the remaining $2-2\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})$ mass is concentrated on binary coalescences, which gives the Kingman part of the $\Xi$-coalescent.
Combining the contributions from $A_{\varepsilon}$ and $A_{\varepsilon}^{\mathrm{c}}$ and setting $q_{n}(\xi,\xi)=-\sum_{\eta\neq\xi}q_{n}(\xi,\eta)$, we conclude that
$c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)\ \to\ Q_{n},$
where $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$. This completes the proof. ∎
Remark Appendix A.2.
[10, Theorem 1.1] establishes that, if Assumption 3.3 and (Appendix A.1) hold, then the rescaled ancestral process converges to the $\Xi$-coalescent with $\Xi=2\Xi^{\prime}\circ\varphi^{-1}$, where $\varphi$ denotes the halving map $\varphi:\Delta\to\Delta$ defined by
@@
Note that this $\Xi$ (which is under no selfing) is different from the $\Xi=2\Xi^{\prime}$ in the annealed convergence in Theorem 3.9. This difference is expected because when $\alpha_{N}\not\to 1$, the annealed limit of the rescaled ancestral process will not have $Q_{n}$ as its generator. The haploid map $F_{\rm hap}$ in (3.10), which appears in the definition of $Q_{n}$, makes $Q_{n}$ the generator of the limiting process where two blocks that enter the same individual always coalesce instantaneously in the limit. When $\alpha_{N}\not\to 1$, blocks of the ancestral process that enter the same individual may disperse before coalescing, and this is captured in the true limiting generator.
We refer the interested reader to [10, Appendix A] for several equivalent formulations to Equation (Appendix A.1).
Appendix Appendix B Proof of the Theorem 3.9
We produce a small lemma here to show that, for our notion of weak convergence in distribution that the intensity measure map sending a random measure $\mu$ to its intensity measure $\mathbb{E}\left[\mu\right]$ is continuous.
Lemma Appendix B.1.
Suppose that $\mu_{N}$ is a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ that converge weakly in distribution, in the sense of Definition 6.1, and that $E$ is a locally compact Polish space. Then the sequence of intensity measures $\mathbb{E}\left[\mu_{N}\right]$ converge weakly in distribution to $\mathbb{E}\left[\mu\right]$ when $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ is given the topology of uniform convergence on compacta.
Proof.
Recall that for a random measure $\mu$ that the intensity measure $\mathbb{E}\left[\mu\right]$ is defined as the unique measure for which
$\mathbb{E}\left[\mu(f)\right]=\mathbb{E}\left[\mu\right](f)$
for any $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
The topology on $\mathcal{M}_{1}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ is such that the evaluation maps $T_{f}$ sending $\mu$ to $\mu(f)$ is continuous and bounded. By the continuous mapping theorem therefore $T_{f}(\mu_{N})=\mu_{N}(f)$ converges in distribution to $T_{f}(\mu)=\mu(f)$. Because $T_{f}$ is continuous and bounded, convergence in distribution implies convergence in expectation. This yields the claim.
With Theorem 4.7 in hand, we are ready to prove Theorem 3.9.
Proof.
Note that the assumptions of Theorem 3.9 are the same as those of Theorem 4.7. Consequently, by Theorem 4.7 we have that the sequence of random measures $\mathcal{L}^{N,n}$ converges weakly in distribution to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is an EFC with coalescence measure $\Xi$ on $\Delta$, no non-binary fragmentation, and binary fragmentation with rate $\lambda$. By Lemma Appendix B.1, the intensity measure mapping is continuous, and so $\mathbb{E}\left[\mathcal{L}^{N,n}\right]=\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\right)$ converges weakly to $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$. It suffices, therefore, to show that $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$ is the law of an $n$-$\Xi$-coalescent.
@@
Let $N$ be fixed and consider the general diploid overlapping-generations model described at the beginning of Section 2. At time step 0, let $K:=K_{N}^{(0)}$, $S:=S_{N}^{(0)}=\sum_{i=1}^{N}V^{(0)}_{i,i}$, $v_{i}:=V^{(0)}_{i,i}$ and $u_{i}:=\sum_{j\neq i}V^{(0)}_{i,j}$ for $i\in[N]$.
@@
$\mathfrak{c}_{3}:=\mathbb{P}_{\xi_{0}^{3}}\left(\chi^{N,n}(1)=\mathbf{1}_{3}\right)=\frac{1}{\binom{N}{3}}\,$ $\mathbb{E}\!\bigg[\frac{N-K}{N}\sum_{i=1}^{N}\!\left(\frac{1}{4}\binom{v_{i}}{2}+\frac{1}{8}\,v_{i}u_{i}+\frac{1}{16}\binom{u_{i}}{2}\right)$
@@
Sample uniformly without replacement three distinct individuals at time 0; there are $\binom{N}{3}$ unordered choices. Condition first on the realized newborn set $B$ (of size $K$), the carried-over set C~C (of size NKN-K), and the parentage matrix $V^{(0)}$.
There are exactly two ways to obtain a one-step triple merger:
(A) Two newborns share a parent $i$ and the third sampled individual is that parent $i$ carried over. Fix $i\in C$. Among the $n_{i}:=v_{i}+u_{i}$ newborns that have $i$ as a parent, there are $\binom{v_{i}}{2}$ pairs where both are selfed by $i$, $v_{i}u_{i}$ mixed pairs, and $\binom{u_{i}}{2}$ pairs with both outcrossing via $i$. Tracing one step back, to coalesce with the gene copy traced from the carried-over $i$, each newborn must choose the same copy in $i$: probabilities $1/2$ (selfed) and $1/4$ (outcrossed), independently of the carried-over copy choice. Hence the Mendelian factors are $1/4$, $1/8$, and $1/16$, respectively. Summing over $i\in C$ gives the conditional contribution
@@
(B) Three newborns share a common parent $i$ (not necessarily carried over). For a fixed $i\in[N]$, the three newborns can be all selfed ($\binom{v_{i}}{3}$ triples), two selfed and one outcrossed ($\binom{v_{i}}{2}u_{i}$ triples), one selfed and two outcrossed ($v_{i}\binom{u_{i}}{2}$ triples), or all outcrossed ($\binom{u_{i}}{3}$ triples). To coalesce one step back, all three lineages must choose the same copy in $i$, yielding Mendelian factors $1/8$, $1/16$, $1/32$, and $1/64$, respectively. Summing over $i$ gives the conditional contribution
@@
Adding (A) and (B) yields the conditional probability of triple coalescent given $V^{(0)},B,C$ as
@@
Suppose that $\mathfrak{c}_{3}\in o(c_{N})$ as $N\to\infty$. Then the one-step transition probability $\mathit{p}_{\xi\eta}^{N,n}$ for any $\eta$ that cannot be obtained from $\xi$ via a single binary merger, i.e. where $\xi\not\prec\eta$ , satisfies $\mathit{p}_{\xi\eta}^{N,n}c_{N}^{-1}\to 0$. Further, $\mathit{p}_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$.
@@
Let $r$ denote the largest number of blocks of $\xi$ that are coalesced together into a single block of $\eta$. Suppose to start that $r\geq 3$ after factoring through the haploid map. For any subset of size $3$ of these $r$ blocks, of which there are $\binom{r}{3}$, the probability of that given subset coalescing in a single time-step is $\mathfrak{c}_{3}$. Therefore the one-step transition probability $\mathit{p}_{\xi\eta}^{N,n}$ that $\eta$ is obtained from $\xi$ is at most $\binom{r}{3}\mathfrak{c}_{3}$, which is $o(c_{N})$ by assumption. This shows that any $\eta$ obtained by coalescing $r\geq 3$ blocks of $\xi$ into a single block happens with $o(c_{N})$ probability.
@@
The second summand is $O(c_{N}^{2})\subset o(c_{N})$ by a direct comparison, and by assumption $\mathfrak{c}_{3}$ is $o(c_{N})$. This shows that when $r=2$ with $\xi\not\prec\eta$ that $\mathit{p}_{\xi\eta}^{N,n}\in o(c_{N})$.

### v2

@@ Acknowledgements
$\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x):=\sum_{l=0}^{s}\sum_{i_{1},\ldots,i_{r}\ \mathrm{distinct}}\binom{s}{l}\Big(\prod_{j=1}^{r}x_{i_{j}}^{k_{j}}\Big)\left(\prod_{j=r+1}^{l}x_{i_{j}}\right)\,(1-|x|)^{s-l}\frac{\Xi(dx)}{\left\langle x,x\right\rangle}$
the probability that $b$ blocks thrown uniformly on $[0,1]$ with a paintbox $x$ form a partition of type $k_{1},\ldots,k_{r};s$ (see [9, Equation 1.9]). For any measure $\Theta$ on $\Delta$, the probability that $b$ blocks thrown uniformly at random and independently on $[0,1]$ have $r$ groups of size $k_{1},\ldots,k_{r}$ falling into the same intervals and $s$ blocks falling into their own unique intervals (or else in excess of $\sum_{i}x_{i}$) for $x$ sampled from $\frac{1}{\left\langle x,x\right\rangle}\Theta$ is exactly
$\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)}{\left\langle x,x\right\rangle}d\Theta(x).$ (Appendix B.6)
That is, the $x$-paintbox transition probability for $x$ sampled from $\frac{1}{\left\langle x,x\right\rangle}\Theta$ is exactly (Appendix B.6).
We now show that
$\lim_{\varepsilon\to 0}\lim_{N\to\infty}\Phi_{N}(A_{\varepsilon})Q_{N,n}^{\varepsilon}(\xi_{0}^{b},\eta)=2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)}{\left\langle x,x\right\rangle}d\Xi^{\prime}(x),$ (Appendix B.7)
which implies that $\lim_{\varepsilon\to 0}\lim_{N\to\infty}\Phi_{N}(A_{\varepsilon})Q_{N,n}^{\varepsilon}$ exists and is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}-2\delta_{\mathbf{0}}\Xi^{\prime}(\mathbf{0})$. Conditional on $x\neq\mathbf{0}$, we have
$\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)=2\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)+\Delta_{N}(x),\qquad\text{with }|\Delta_{N}(x)|\leq\frac{K_{b}}{N}.$
That $|\Delta_{N}(x)|$ is $O(\frac{1}{N})$ follows simply from the observation that the probability of the $n\ll N\sum_{i}x_{i}$ sample lineages falling into each of the intervals of $x$ are asymptotically independent, with corrections at worst of order $\frac{1}{N}$, which can be uniformly controlled by compactness of $\Delta$ and finiteness of $b$. Therefore,
$\frac{1}{c_{N}}\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)\right]=\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,2\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)\right]+o(1),$
since the $\frac{1}{N}$ coupling error is multiplied by $\Phi_{N}(A_{\varepsilon})/c_{N}=O_{\varepsilon}(1)$. By Assumption (Appendix B.4), vague convergence on $\Delta\setminus\{\mathbf{0}\}$ yields
$\lim_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)\right]=2\int_{A_{\varepsilon}}\frac{\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
@@
$2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
which coincides with the $x$-paintbox transition probability for $\Xi\!\restriction_{\Delta\setminus\{\mathbf{0}\}}=2\Xi^{\prime}\!\restriction_{\Delta\setminus\{\mathbf{0}\}}$.
We proceed to show now that
$\lim_{\varepsilon\to 0}\lim_{N\to\infty}Q_{N,n}^{\epsilon,c}(\xi_{0}^{b},\eta)=\begin{cases}0&,\xi\not\prec\eta\\ 2&,\xi\prec\eta,\end{cases},$ (Appendix B.8)
where $\xi\prec\eta$ if $\eta$ is the result of a binary merger of blocks in $\xi$. That is, we will show $\lim_{\varepsilon\to 0}\lim_{N\to\infty}Q_{N,n}^{\epsilon,c}$ exists and is equal to the generator of an $n$-$\Xi$-coalescent with $\Xi=\Xi^{\prime}(\mathbf{0})\delta_{\mathbf{0}}$. It suffices, therefore, to show that non-binary mergers occur with negligible probability and that binary mergers occur with rate $2$. For $x\in A_{\varepsilon}^{\mathrm{c}}$ we have $\langle x,x\rangle\leq\varepsilon$, hence $\max_{i}x_{i}\leq\|x\|_{2}\leq\varepsilon^{1/2}$. Then
@@
Let $M_{b}$ denote the event that, among the $b$ blocks thrown on $[0,1]$ that there is at least one non-binary merger from the $x$-paintbox. Then
$\mathbb{P}(M_{b}\mid x)\;\leq\;C_{b}\!\left(\sum_{i}x_{i}^{3}+\Big(\sum_{i}x_{i}^{2}\Big)^{2}\right)\;\leq\;C_{b}\big(\varepsilon^{1/2}+\varepsilon\big)\sum_{i}x_{i}^{2}\;\leq\;2C_{b}\,\varepsilon^{1/2}\sum_{i}x_{i}^{2}$
for some constant $C_{b}$ depending only on $b$. Therefore, using $\lim_{N\to\infty}\frac{1}{2c_{N}}\mathbb{E}\!\big[\sum_{i}x_{i}^{2}\big]=\Xi^{\prime}(\Delta\setminus{\{\mathbf{0}}\})$,
$\limsup_{N\to\infty}\frac{1}{2c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(M_{b}\mid x)\big]\;\leq\;2C_{b}\,\varepsilon^{1/2}\,\limsup_{N\to\infty}\frac{1}{2c_{N}}\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]\;=\;2C_{b}\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})\,\varepsilon^{1/2}.$
@@
$\lim_{\varepsilon\downarrow 0}\;\limsup_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(M_{b}\mid x)\big]=0.$
@@
and on $A_{\varepsilon}^{\mathrm{c}}$ the $O(\sum_{i}x_{i}^{3})$ error satisfies
@@
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]=2-\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}}\sum_{i}x_{i}^{2}\big].$
By the normalization and the vague convergence assumption (Appendix B.4), as $N\to\infty$,
@@
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]\;\to\;2-2\Xi^{\prime}(A_{\varepsilon})\xrightarrow[\varepsilon\downarrow 0]{}\;2\Xi^{\prime}(\mathbf{0}).$
Combining the contributions from $A_{\varepsilon}$ from (Appendix B.7) and $A_{\varepsilon}^{\mathrm{c}}$ from (Appendix B.8) with the decomposition (Appendix B.5) yields
$\lim_{N\to\infty}Q_{N,n}(\xi,\eta)$ $=\lim_{\varepsilon\to 0}\lim_{N\to\infty}\Phi_{N}(A_{\varepsilon})\,Q_{N,n}^{\varepsilon}+\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\,Q_{N,n}^{\varepsilon,\mathrm{c}}$
$=\begin{cases}2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Upsilon_{b;(2;b-2)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx)+2\Xi^{\prime}(\mathbf{0})&,\text{ if }\xi\prec\eta\\ 2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Upsilon_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx)&,\text{ if }\xi\not\prec\eta\end{cases}.$
This is precisely the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$ by comparing with (Appendix B.2).
Remark Appendix B.3.
[9, Theorem 1.1] establishes that, if Assumption 3.4 and (Appendix B.4) hold, then the rescaled ancestral process converges to the $\Xi$-coalescent governed by $\Xi^{\prime}\circ\varphi^{-1}$, where $\varphi$ denotes the halving map $\varphi:\Delta\to\Delta$ defined by
@@
Note that this measure $\Xi^{\prime}\circ\varphi^{-1}$ (which is under no selfing) is different from the measure $2\Xi^{\prime}$ in the annealed convergence in Theorem 3.11. This difference is expected because when $\alpha_{N}\not\to 1$, blocks of the ancestral process that enter the same individual may disperse before coalescing, and the annealed limit of the rescaled ancestral process will not have $Q_{n}$ as its generator. The reason is that the haploid map $F_{\rm hap}$ in (3.9), which appears in the definition of $Q_{n}$, makes $Q_{n}$ the generator of the limiting process where two blocks that enter the same individual always coalesce instantaneously in the limit.
We refer the interested reader to [9, Appendix A] for several equivalent formulations to Equation (Appendix B.4).
@@
Let $N$ be fixed and consider the general diploid overlapping-generations model described at the beginning of Section 2. At time step 0, let $K:=K_{N}^{(0)}$, $S:=\sum_{i=1}^{N}V^{(0)}_{i,i}$, $v_{i}:=V^{(0)}_{i,i}$ and $u_{i}:=\sum_{j\neq i}V^{(0)}_{i,j}$ for $i\in[N]$.
@@
$\mathfrak{c}_{3}=\frac{1}{\binom{N}{3}}\,$ $\mathbb{E}\!\bigg[\frac{N-K}{N}\sum_{i=1}^{N}\!\left(\frac{1}{4}\binom{v_{i}}{2}+\frac{1}{8}\,v_{i}u_{i}+\frac{1}{16}\binom{u_{i}}{2}\right)$
@@
Sample uniformly without replacement three distinct individuals at time 0; there are $\binom{N}{3}$ unordered choices. Condition first on the realized newborn set $B$ (of size $K$), the carried-over set C~C (of size NKN-K), and the parentage matrix $V^{(0)}$. There are two ways to obtain a one-step triple merger:
(i) Two newborns share a parent $i$ and the third sampled individual is that parent $i$ carried over. Fix $i\in C$. Among the $n_{i}:=v_{i}+u_{i}$ newborns that have $i$ as a parent, there are $\binom{v_{i}}{2}$ pairs where both are selfed by $i$, $v_{i}u_{i}$ mixed pairs, and $\binom{u_{i}}{2}$ pairs with both outcrossing via $i$. Tracing one step back, to coalesce with the gene copy traced from the carried-over $i$, each newborn must choose the same copy in $i$: probabilities $1/2$ (selfed) and $1/4$ (outcrossed), independently of the carried-over copy choice. Hence the Mendelian factors are $1/4$, $1/8$, and $1/16$, respectively. Summing over $i\in C$ gives the conditional contribution
@@
(ii) Three newborns share a common parent $i$ (not necessarily carried over). For a fixed $i\in[N]$, the three newborns can be all selfed ($\binom{v_{i}}{3}$ triples), two selfed and one outcrossed ($\binom{v_{i}}{2}u_{i}$ triples), one selfed and two outcrossed ($v_{i}\binom{u_{i}}{2}$ triples), or all outcrossed ($\binom{u_{i}}{3}$ triples). To coalesce one step back, all three lineages must choose the same copy in $i$, yielding Mendelian factors $1/8$, $1/16$, $1/32$, and $1/64$, respectively. Summing over $i$ gives the conditional contribution
@@
Adding the above two displays yields the conditional probability of triple coalescent given $V^{(0)},B,C$ as
@@
Suppose that $\mathfrak{c}_{3}\in o(c_{N})$ as $N\to\infty$. Then for any $\xi\in\mathcal{E}_{n}$, it holds that $\mathit{h}_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$ and that $\mathit{h}_{\xi\eta}^{N,n}\in o(c_{N})$ for any $\eta$ that cannot be obtained from $\xi$ via a single binary merger, i.e. where $\xi\not\prec\eta$
@@
Let $r$ denote the largest number of blocks of $\xi$ that are coalesced together into a single block of $\eta$. Suppose to start that $r\geq 3$ after factoring through the haploid map. For any subset of size $3$ of these $r$ blocks, of which there are $\binom{r}{3}$, the probability of that given subset coalescing in a single time-step is $\mathfrak{c}_{3}$. Therefore the one-step transition probability $h_{\xi\eta}^{N,n}$ that $\eta$ is obtained from $\xi$ is at most $\binom{r}{3}\mathfrak{c}_{3}$, which is $o(c_{N})$ by assumption. This shows that any $\eta$ obtained by coalescing $r\geq 3$ blocks of $\xi$ into a single block happens with $o(c_{N})$ probability.
@@
The second summand is $O(c_{N}^{2})\subset o(c_{N})$ by a direct comparison, and by assumption $\mathfrak{c}_{3}$ is $o(c_{N})$. This shows that when $r=2$ with $\xi\not\prec\eta$ that $h_{\xi\eta}^{N,n}\in o(c_{N})$.